# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ Appendix A Structured Representation of Nonlinear Functions
In Section II, we have reviewed two popular frameworks for representing (nonlinear) functions: reproducing kernel Hilbert spaces (RKHS) and neural network function spaces (NNFS). Typical kernel functions $\ker(\cdot,\cdot)$ to define RKHSs include
  1. 1.
    Gaussian kernel,
  2. 2.
    Matern kernel,
  3. 3.
    Linear kernel,
  4. 4.
    Laplacian kernel,
  5. 5.
    Polynomial kernel.
Mathematical details of these kernel functions can be found in [24, Subsec. 14.2], [27, Ex. 1]. Typical activation functions $\sigma(\cdot)$ to define NNFSs include
  1. 1.
    Hyperbolic tangent (i.e, tanh) function,
  2. 2.
    Softmax function,
  3. 3.
    Sigmoid function,
  4. 4.
    Rectified linear unit (ReLU) function,
  5. 5.
    Exponential linear unit (ELU) function,
  6. 6.
    Linear function.
Mathematical details of these activation functions can be found in [27, Ex. 2].
@@ Appendix B Details on Real-Space Signal Representation
We have
@@
$\begin{array}[]{cl}\bm{R}_{\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{s}}\underline{\mathbf{s}}^{\mathsf{T}}}&=\displaystyle\frac{1}{2}\left[\begin{array}[]{cc}\operatorname{Re}(\bm{P}+\bm{C})&\operatorname{Im}(-\bm{P}+\bm{C})\\ \operatorname{Im}(\bm{P}+\bm{C})&\operatorname{Re}(\bm{P}-\bm{C})\end{array}\right],\end{array}$
@@
$\begin{array}[]{cl}\bm{R}_{\underline{v}}\coloneqq\mathbb{E}{\underline{\mathbf{v}}\underline{\mathbf{v}}^{\mathsf{T}}}&=\displaystyle\frac{1}{2}\left[\begin{array}[]{cc}\operatorname{Re}\bm{R}&\operatorname{Im}-\bm{R}\\ \operatorname{Im}\bm{R}&\operatorname{Re}\bm{R}\end{array}\right].\end{array}$
Note that notationally $\bm{P}\coloneqq\bm{R}_{s}$, $\bm{C}\coloneqq\bm{C}_{s}$, and $\bm{R}\coloneqq\bm{R}_{v}$. Note also that the following identities hold: $\bm{R}_{x}=\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}$, $\bm{C}_{x}=\bm{H}\bm{C}_{s}\bm{H}^{\mathsf{T}}$, $\bm{R}_{\underline{x}}=\underline{\underline{\bm{H}}}\cdot\bm{R}_{\underline{s}}\cdot\underline{\underline{\bm{H}}}^{\mathsf{T}}+\bm{R}_{\underline{v}}$, and $\bm{R}_{\underline{xs}}=\underline{\underline{\bm{H}}}\cdot\bm{R}_{\underline{s}}$.
@@ Appendix C Extensive Reading on Distributional Uncertainty
$\begin{array}[]{l}\displaystyle\min_{\bm{\phi}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}\\ \quad\quad=\displaystyle\min_{\bm{\phi}}\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}\\ \quad\quad=\displaystyle\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}^{\star}_{\text{ERM}}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}^{\star}_{\text{ERM}}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}\\ \quad\quad\leq\displaystyle\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}^{\star}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}^{\star}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}.\end{array}$ (62)
@@
In the practice of machine learning, we want to reduce both the generalization error and the excess risk. Most attention in the literature has been particularly paid to controlling generalization errors. To be specific, an upper bound of the true cost $\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\mathbb{P}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$ is first found and then minimize the upper bound: by minimizing the upper bound, the true cost can also be reduced. For technical details, see Fact 1.
Appendix D Wasserstein Distance and Maximum Mean Discrepancy Distance
Consider two distributions (probability measures) $\mathbb{P}_{\mathbf{x}}$ and $\mathbb{P}_{\mathbf{y}}$ on the same measurable space $(\Xi,\mathcal{B}_{\Xi})$ where $\mathcal{B}_{\Xi}$ is the Borel $\sigma$-algebra on the set $\Xi$. The Wasserstein distance $d_{\text{W}}$ between $\mathbb{P}_{\mathbf{x}}$ and $\mathbb{P}_{\mathbf{y}}$ is defined as [27, Def. 2]
$d_{\text{W}}(\mathbb{P}_{\mathbf{x}},\mathbb{P}_{\mathbf{y}})\coloneqq\inf_{\Pi}\int_{\Xi\times\Xi}d(\bm{x},\bm{y})\Pi(\mathrm{d}\bm{x},\mathrm{d}\bm{y}),$
where $d$ is a metric on $\Xi$ and $\Pi$ is a joint distribution of $\mathbf{x}$ and $\mathbf{y}$; $\Xi$ is a coupling of $\mathbb{P}_{\mathbf{x}}$ and $\mathbb{P}_{\mathbf{y}}$: i.e., the marginal distributions induced by $\Pi$ is $\mathbb{P}_{\mathbf{x}}$ and $\mathbb{P}_{\mathbf{y}}$.
The maximum mean discrepancy (MMD) distance $d_{\text{MMD}}$ between $\mathbb{P}_{\mathbf{x}}$ and $\mathbb{P}_{\mathbf{y}}$ in the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ is defined as [28, Def. 2.1]
$d_{\text{MMD}}(\mathbb{P}_{\mathbf{x}},\mathbb{P}_{\mathbf{y}})\coloneqq\sup_{f\in\mathcal{H}:~{}\|f\|_{\mathcal{H}}\leq 1}\left[\mathbb{E}_{\mathbf{x}}f(\bm{x})-\mathbb{E}_{\mathbf{y}}f(\bm{y})\right],$
where $\|\cdot\|_{\mathcal{H}}$ is a norm on $\mathcal{H}$ [24, p. 539].
@@ Appendix E Proof of Theorem 1
Since $\bm{R}_{x,1}\succeq\bm{R}_{x,2}$, we have
$\begin{array}[]{l}\left[\begin{array}[]{cc}\bm{R}_{s}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x,1}\end{array}\right]-\left[\begin{array}[]{cc}\bm{R}_{s}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x,2}\end{array}\right]\\ \quad\quad=\left[\begin{array}[]{cc}\bm{0}&\bm{0}\\ \bm{0}&\bm{R}_{x,1}-\bm{R}_{x,2}\end{array}\right]\succeq\bm{0}.\end{array}$
According to the properties of the Schur complement, we have $f(\bm{R}_{x,1})\geq f(\bm{R}_{x,2})$. $\square$

### v2

@@ Appendix A Structured Representation of Nonlinear Functions
In Section II, we have reviewed two popular frameworks for representing (nonlinear) functions: reproducing kernel Hilbert spaces (RKHS) and neural network function spaces (NNFS). Typical kernel functions $\ker(\cdot,\cdot)$ to define RKHSs include Gaussian kernel, Matern kernel, Linear kernel, Laplacian kernel, and Polynomial kernel. Mathematical details of these kernel functions can be found in [24, Subsec. 14.2], [27, Ex. 1]. Typical activation functions $\sigma(\cdot)$ to define NNFSs include Hyperbolic tangent (i.e, tanh) function, Softmax function, Sigmoid function, Rectified linear unit (ReLU) function, and Exponential linear unit (ELU) function. Mathematical details of these activation functions can be found in [27, Ex. 2].
@@ Appendix B Details on Real-Space Signal Representation
Let $\bm{C}_{x}\coloneqq\mathbb{E}\mathbf{x}\mathbf{x}^{\mathsf{T}}$, $\bm{C}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{T}}$, and $\bm{C}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{T}}=\bm{0}$. We have
@@
$\begin{array}[]{cl}\bm{R}_{\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{s}}\underline{\mathbf{s}}^{\mathsf{T}}}&=\displaystyle\frac{1}{2}\left[\begin{array}[]{cc}\operatorname{Re}(\bm{R}_{s}+\bm{C}_{s})&\operatorname{Im}(-\bm{R}_{s}+\bm{C}_{s})\\ \operatorname{Im}(\bm{R}_{s}+\bm{C}_{s})&\operatorname{Re}(\bm{R}_{s}-\bm{C}_{s})\end{array}\right],\end{array}$
@@
$\begin{array}[]{cl}\bm{R}_{\underline{v}}\coloneqq\mathbb{E}{\underline{\mathbf{v}}\underline{\mathbf{v}}^{\mathsf{T}}}&=\displaystyle\frac{1}{2}\left[\begin{array}[]{cc}\operatorname{Re}\bm{R}_{v}&\operatorname{Im}-\bm{R}_{v}\\ \operatorname{Im}\bm{R}_{v}&\operatorname{Re}\bm{R}_{v}\end{array}\right].\end{array}$
Note that the following identities hold: $\bm{R}_{x}=\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}$, $\bm{C}_{x}=\bm{H}\bm{C}_{s}\bm{H}^{\mathsf{T}}$, $\bm{R}_{\underline{x}}=\underline{\underline{\bm{H}}}\cdot\bm{R}_{\underline{s}}\cdot\underline{\underline{\bm{H}}}^{\mathsf{T}}+\bm{R}_{\underline{v}}$, and $\bm{R}_{\underline{xs}}=\underline{\underline{\bm{H}}}\cdot\bm{R}_{\underline{s}}$.
@@ C-A Generalization Error and Distributional Robustness
$\begin{array}[]{l}\displaystyle\min_{\bm{\phi}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}\\ \quad\quad=\min_{\bm{\phi}}\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}\\ \quad\quad=\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}^{\star}_{\text{ERM}}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}^{\star}_{\text{ERM}}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}\\ \quad\quad\leq\operatorname{Tr}\frac{1}{L}\sum^{L}_{i=1}[\bm{\phi}^{\star}(\bm{x}_{i})-\bm{s}_{i}][\bm{\phi}^{\star}(\bm{x}_{i})-\bm{s}_{i}]^{\mathsf{H}}.\end{array}$ (73)
@@
is called the excess risk of $\bm{\phi}^{\star}_{\text{ERM}}$. In machine learning practice, we want to reduce both the generalization error and the excess risk. Most attention in the literature has been particularly paid to reducing generalization errors. Specifically, an upper bound of the true cost $\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\mathbb{P}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$ is first found and then minimize the upper bound: by minimizing the upper bound, the true cost can also be reduced.
Fact 1
Suppose that the true distribution $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ of $(\mathbf{x},\mathbf{s})$ is included in $\mathcal{U}_{\mathbf{x},\mathbf{s}}$; for notational clarity, we hereafter distinguish $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ from $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. The true objective function evaluated at $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$, i.e.,
$\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\mathbb{P}_{0,\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}},~{}~{}~{}\forall\bm{\phi}\in\mathcal{B},$ (74)
is upper bounded by the worst-case objective function of (17), i.e.,
$\max_{\mathbb{P}_{\mathbf{x},\mathbf{s}}\in\mathcal{U}_{\mathbf{x},\mathbf{s}}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\mathbb{P}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}},~{}~{}~{}\forall\bm{\phi}\in\mathcal{B}.$ (75)
Therefore, by diminishing the upper bound in (75), the true estimation error evaluated at $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ can also be reduced. However, the conventional empirical estimation error evaluated at ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ cannot upper bound the true estimation error (74). This performance guarantee is the benefit of considering the distributionally robust method (17). Due to the weak convergence property of the empirical distribution to the true data-generating distribution, that is, $d(\mathbb{P}_{0,\mathbf{x},\mathbf{s}},~{}\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}})\to 0$ as the sample size $L\to\infty$, there exists $\epsilon$ in (18) for every $L$, such that $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ is included in $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ in $\mathbb{P}^{L}_{0,\mathbf{x},\mathbf{s}}$-probability ($L$-fold product measure of $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$). $\square$
C-B Non-Stationary Channel Statistics
In the main body of the paper (see also Fact 1), we assume that the true data-generating distribution $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ is time-invariant within a frame. In real-world operations, however, this assumption might be untenable.
As shown in Fig. 1, the frame contains eight data units; we suppose that the first four units are pilot symbols and the rest four units are communication-data symbols.
Figure 1: True data-generating distributions might be time-varying in a frame.
Let $\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ denote the true data-generating distribution at time point $t_{i}$ where $i=0,1,2,\ldots,8$. Specifically, we have $(\mathbf{x}_{i},\mathbf{s}_{i})\sim\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ for every $i$. Therefore, the pilot data set (i.e., the training data set) $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),(\bm{x}_{3},\bm{s}_{3}),(\bm{x}_{4},\bm{s}_{4})\}$ can be seen as realizations of the mean distribution $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}$ of underlying true training-data distributions where $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\sum^{4}_{i=1}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, which is a mixture distribution with mixing weights $0\leq h_{1},h_{2},h_{3},h_{4}\leq 1$; $\sum^{4}_{i=1}h_{i}=1$. Similarly, the communication data set (i.e., the testing data set) $\{(\bm{x}_{5},\bm{s}_{5}),(\bm{x}_{6},\bm{s}_{6}),(\bm{x}_{7},\bm{s}_{7}),(\bm{x}_{8},\bm{s}_{8})\}$ can be seen as realizations of the mean $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$ of the underlying true testing-data distributions where $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}=\sum^{8}_{i=5}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, with mixing weights $0\leq h_{5},h_{6},h_{7},h_{8}\leq 1$; $\sum^{8}_{i=5}h_{i}=1$.
Suppose that
$d(\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}},~{}\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}})\leq\epsilon_{1},$
where ^$\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}}\coloneqq\frac{1}{4}\sum^{4}_{i=1}\delta_{(\bm{x}_{i},\bm{s}_{i})}$ is the data-driven estimate of $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}$ and
$d(\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}},~{}\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}})\leq\epsilon_{2},$
for some $\epsilon_{1},\epsilon_{2}\geq 0$. We have the uncertainty quantification
$d(\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}},~{}\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}})\leq\epsilon\coloneqq\epsilon_{1}+\epsilon_{2}.$
Therefore, the distributionally robust modeling and solution framework is still valid to hedge against the distributional uncertainty in the nominal distribution ^$\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}}$ compared to the underlying true distribution $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$. When $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, as assumed in the main body of the paper, we have $\epsilon_{1}\to 0$ and $\epsilon\to\epsilon_{2}=0$ as the pilot size tends to infinity; however, when $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}\neq\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, the radius $\epsilon\to\epsilon_{2}\neq 0$ although $\epsilon_{1}\to 0$.
Another justification for the DRO method is as follows. Suppose that there exists $\epsilon\geq 0$ such that
$d(\mathbb{P}_{0,\mathbf{x},\mathbf{s},i},~{}\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}})\leq\epsilon,~{}~{}~{}\forall i\in\{0,1,2,\ldots,8\}.$
It means that, at every snapshot in the frame, the true data-generating distribution is included in the uncertainty set. Hence, the DRO cost can still upper bound the true cost even though the true distribution is time-varying; cf. Fact 1.
Appendix D Proof of Lemma 1
Proof:
The objective function of Problem (19) equals to
$\left\langle\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right],~{}\left[\begin{array}[]{cc}\bm{R}_{x}&\bm{R}_{xs}\\ \bm{R}^{\mathsf{H}}_{xs}&\bm{R}_{s}\end{array}\right]\right\rangle,$ (76)
where $\langle\bm{A},\bm{B}\rangle\coloneqq\operatorname{Tr}\bm{A}^{\mathsf{H}}\bm{B}$ for two matrices $\bm{A}$ and $\bm{B}$. Therefore, the objective function of (19) is convex in $\bm{W}$ and linear (thus concave) in the matrix variable $\bm{R}$. Note also that complex Euclidean spaces are linear topological spaces. Hence, due to Sion’s minimax theorem, Problem (19) is equivalent to
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}\min_{\bm{W}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\bm{R}_{xs}-\bm{R}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0},\\ &\bm{R}\succeq\bm{0}.\end{array}$ (77)
For every given $\bm{R}$, the inner minimization sub-problem of (77) is solved by the Wiener beamformer $\bm{W}^{\star}_{\text{Wiener}}=\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}$, which transforms (77) to (21). This completes the proof. $\square$
@@ Appendix E Proof of Theorem 1
Consider the following optimization problem
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\bm{R}\succeq\bm{R}_{2},\\ &\bm{R}_{x}\succ\bm{0},\end{array}$ (78)
which, due to Lemma 1, is equivalent [in the sense of the same optimal objective value and maximizer(s) $\bm{R}^{\star}$] to
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}}&\left\langle\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right],~{}\left[\begin{array}[]{cc}\bm{R}_{x}&\bm{R}_{xs}\\ \bm{R}^{\mathsf{H}}_{xs}&\bm{R}_{s}\end{array}\right]\right\rangle\\ \text{s.t.}&\bm{R}\succeq\bm{R}_{2},\\ &\bm{R}_{x}\succ\bm{0}.\end{array}$ (79)
Note that $\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right]\succeq\bm{0},$ because for all $\bm{x}\in\mathbb{C}^{N}$ and $\bm{y}\in\mathbb{C}^{M}$, we have
$[\bm{x}^{\mathsf{H}},~{}\bm{y}^{\mathsf{H}}]\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right]\left[\begin{array}[]{c}\bm{x}\\ \bm{y}\end{array}\right]=\|\bm{W}\bm{x}-\bm{y}\|^{2}_{F}\geq 0.$
Therefore, for every given $\bm{W}$, the objective function of (79) is increasing in $\bm{R}$. As a result, the objective value of (78) is lower-bounded at $\bm{R}_{2}$: To be specific, $\forall\bm{R}\succeq\bm{R}_{2}$, we have
$\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \geq\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{2,xs}\bm{R}^{-1}_{2,x}\bm{R}_{2,xs}+\bm{R}_{2,s}\big{]},$
i.e., $f_{1}(\bm{R})\geq f_{1}(\bm{R}_{2})$, which proves the first part.
On the other hand, if $\bm{R}_{1,x}\succeq\bm{R}_{2,x}\succ\bm{0}$, we have $\bm{R}^{-1}_{2,x}\succeq\bm{R}^{-1}_{1,x}$. As a result, $f_{2}(\bm{R}_{1,x})-f_{2}(\bm{R}_{2,x})=\operatorname{Tr}\left[{\bm{R}^{\mathsf{H}}_{xs}(\bm{R}^{-1}_{2,x}-\bm{R}^{-1}_{1,x})\bm{R}_{xs}}\right]\geq 0$, completing the proof. $\square$